# Bridging Gulfs in UI Generation through Semantic Guidance

### v1

@@ 5.2. System Overview
As shown in Figure 3, users first specify semantics in the semantic structure view across the four-level framework (top-left); the system then generates the corresponding UI from these semantics and renders it (top-right). This pairing of semantic input and visual feedback anchors user intent to concrete outcomes, reducing ambiguity. Users can change to the relation view (bottom-left), which visualizes interdependencies among semantics. From either view, clicking any semantic attribute opens the semantic detail panel for in-depth inspection of that attribute and its relationships.
@@ 5.4. Augmented Semantics Analysis
5.5. Semantic Relationship Analysis
As shown in Figure 6, the relationship analysis renders a relation graph between semantic attributes. The graph represents semantics as nodes and their relationships as edges, with three relationship types: Values Match Well [Uncaptioned image] in green, Values Conflict [Uncaptioned image] in orange, and Needs Value [Uncaptioned image] in blue dashed line. This graph serves dual purposes: guiding input decisions before generation and explaining implementation patterns after generation.
@@ 5.6. Iterative Refinement Support
First, the system preserves the semantic state across iterations and utilizes it in every generation. Second, to support granular, section-specific modifications, we designed a targeted regeneration strategy. While standard chat interfaces rely on the model to implicitly infer changes from a stream of conversation history, our system explicitly specifies the semantic diff (comparing previous semantics and identifying which semantic value changed and how) into the prompt. Given that the model regenerates the entire component, by providing previously generated UI code as a reference, the system instructs the model to apply changes to the corresponding modified semantics while preserving the existing structure. This strategy enables a controlled, section-specific update, preventing unintended side effects even when regenerating the full component.
@@ 6.1. Participants and Study Design
We recruited UI/UX practitioners through online design communities, resulting in 14 participants (8 UI/UX designers, 3 developers, 3 product managers) with varying levels of experience in UI generation tools (6 beginners, 5 intermediate users, 3 experts). Participants had an average of 4 years of professional experience (SD = 2.48, range: 1-10 years). Participants received compensation equivalent to US\$20 upon completing the session.
We employed a within-subjects design where each participant used both systems: our semantic-based interface and a baseline chat-based system with v0’s prompt enhancement. Both systems used the identical UI generation model (v0-1.5-md) to ensure that differences stemmed from the interface design rather than generation capabilities.
@@
  • Task 2 (40 mins): Open-ended screen generation with both systems, each with a different scenario.
@@
For each task, participants first generated an initial UI, then were encouraged to refine the generated UI based on their own design goals. To ensure sufficient engagement with the iterative capabilities of both systems, participants were instructed to perform a minimum of three iteration cycles per task. Beyond this requirement, they were free to continue iterating until they reached a state of subjective satisfaction. This structure allowed us to observe both gulfs and amplification problems.
@@ 6.3.1. Quantitative Results
Given the paired design and Likert scale nature of our data, we employed Wilcoxon signed-rank tests (Wilcoxon, 1945) to analyze differences between systems. The tests revealed statistically significant differences across all measured dimensions (*$p<.05$, **$p<.01$) with large effect sizes ($r=z/\sqrt{n}>.5$). Figure 7 illustrates these differences through distinct response patterns. The baseline chat interface produced polarized outcomes—users either succeeded or struggled significantly, with little middle ground. In contrast, our semantic system supported more consistent outcomes, with responses clustering toward positive ratings.
@@
Participants found our system significantly more effective for expressing design intent. They rated Intent Expressiveness substantially higher ($p$ = .008, $r$ = 0.79), with semantic system ratings (M = 5.93) notably exceeding the baseline (M = 4.64). The diverging response patterns in Figure 7 show that while baseline users frequently selected neutral or negative ratings, semantic system users concentrated on positive ratings (6-7 on the Likert scale). Intent Reflection—how accurately users felt the system understood their intentions—also showed significant improvement ($p$ = .010, $r$ = 0.71; Semantic: M = 5.79, Baseline: M = 4.36). The Think through metric demonstrated the largest effect ($p$ = .002, $r$ = 0.87), jumping from M = 3.93 in baseline to M = 6.14 in our system, suggesting that participants perceived the structured semantic input as helpful for systematically considering design decisions rather than struggling with prompt formulation.
@@ 6.3.2. Qualitative Results
Regarding parsing accuracy, twelve participants generally found the natural language to semantic conversion accurate and helpful. P5 characterized the parsing function as “organizing my thinking for me,” while P11 was “surprised by how appropriate the categorization was.” Eleven participants also appreciated how the system flagged empty fields in relationship analysis (Needs Value), guiding them to consider overlooked or missing details. P4 described how this feature guided decision-making: “I knew I wanted a brighter design but wasn’t sure exact one. The relation analysis recommended colors based on my ‘Target User’ and ‘Goal,’ so I applied it immediately.” P7 noted that “empty fields and suggested value showed me what was missing from my prompt,” enabling iterative completion.
@@
Regarding semantic-to-UI adherence, participants reported that specified values were generally well-reflected. P2 noted that “since the system tells me what values to input, I could specify clearly, and it felt well-reflected,” while P10 observed that abstract specifications like “warm pastel tones” resulted in coherent choices across related elements and generated UI. However, adherence varied with input specificity—“clearly specified requests came out as expected, while vague requests produced vague results” [P14]—and design novelty, with familiar UI patterns yielding higher fidelity than novel concepts [P6]. The augmenting semantics (analyzing implementation) received positive feedback for extraction accuracy. P8 reported that “the analyzed results were quite accurate” and P4 surfaced “interaction details I would otherwise forget,” and P5 found augmented semantics helpful for discovering implicit decisions, easy to map design intent to which part of the UI. For semantic relationship analysis, conflict and match value indicators were perceived as actionable and reliable in understanding generated UI. P11 remarked that the “red conflict lines” acted as an immediate feedback loop, “telling me what to fix first” without needing to guess why the UI looked wrong. P10 added that seeing the connections helped explain design choices. Card-level feedback and one-click acceptance/replacement were frequently praised for reducing evaluation effort: “I can see which intent maps to which part of the screen” [P3]. Several participants linked this transparency to greater trust and faster decisions.
@@
The most distinct advantage of the semantic system was its ability to support iterative refinement, reducing “semantic drift”—the tendency for AI models to unexpectedly alter unrelated elements during updates. In the baseline chat condition, five participants (P4, P8, P11, P13, P14) explicitly described “ripple effect” failures. P14 noted: “I asked to change the top section… but it randomly changed the button colors at the bottom too.” P8 echoed this, stating that “Chat updates often reset previous design choices.” Nine participants highlighted the semantic structure’s ability to enable “scoped edits,” allowing users to lock in specific decisions. P2 and P9 highlighted the accuracy of these targeted updates, noting the ability to “target just the node I want to change,” ensuring that the rest of the interface remained stable. The accept/replace workflow of augmented semantics and relationships was described as “clickable evolution” that sped iteration [P5, P8]. In contrast, several noted that chat-based refinement sometimes drifted, altering unrelated visual attributes or layout despite narrow requests [P4, P8, P11, P13, P14]. Participants requested lightweight change auditing—“show me exactly what changed and why”—and simple rollback/version snapshots [P1, P8, P12]. They framed these as additive features atop an already predictable loop.
@@
Across roles, participants saw the semantic approach as especially well-suited to early product exploration, structured refinement, and cross-functional collaboration. Non-designers highlighted gains in expressing and checking intent; designers and engineers pointed to component-level semantics as helpful for handoff and consistency [P9, P11, P12, P13]. Several participants wanted a bridge to light direct manipulation to elements in generated UI that writes back to semantics [P4, P7, P8, P10].
Challenges and Error Responses
Participants identified several limitations in our system. Initial confusion around semantic category boundaries was indicated—P2, P6, and P7 reported difficulty distinguishing Design Style from Visual Mood, or Function from Information Architecture. These concerns typically subsided after one or two generation cycles as participants developed intuitions for the categories. For the semantic parsing, some participants noted that long natural-language descriptions sometimes “thinned out” when compressed into semantic slots, suggesting nuance loss in parsing. Regarding relationship analysis, P1 expressed skepticism about trustworthiness (“Where’s the evidence or source?”), while P11 observed occasional cases where the system “focuses on specific intent, not for the whole intent.” Several participants [P3, P5, P7, P11] noted that both system excels at conventional, minimal styles but struggle with highly branded or stylistically distinctive designs. These findings indicate areas for improvement in semantic granularity, parser fidelity, and support for creative divergence.
@@ 7.2. Design Implications
Designers often want to nudge spacing, alignment, or emphasis directly on canvas without losing traceability. Provide light-weight direct manipulation—handles for spacing/alignment, sliders for radius/elevation, swatch pickers—that update the bound slot and record a semantic diff. Offer “Lock to Design System” so tweaks snap to tokenized ranges. This bridge satisfies craft instincts while maintaining explainability and repeatability; start with a small set of high-impact properties and expand after stability testing.
@@ 7.3. Limitations and Future Work
Our investigation has several limitations. First, the semantic vocabulary introduces an initial learning curve; newcomers must internalize terms and relations. Progressive disclosure, examples-as-templates, and automatic slot suggestions may alleviate this burden. Second, slotting long-form intents risks nuance compression. Dual views that co-display the original text and its structured interpretation, along with preservable “verbatim” spans for critical phrasing, can mitigate loss. Third, the breadth of the evaluation remains limited: our evaluation mainly focused on perceived control, reflecting our theoretical focus on the gulfs as cognitive gaps in the user’s mental model. Consequently, the artifact-based difference, as well as longitudinal and team-level co-operative conditions, remains to be empirically validated. Additionally, technical constraints shaped the current implementation. Our system covers a single React component, which helps reduce integration complexity but limits applicability to complex UI architectures. Also, hallucination (including syntax errors in UI code) and unreliable localized editing remain challenges: instead of modifying only the intended region, the model may regenerate or alter unaffected content. Semantic diffs helped constrain these behaviors, but they serve as a temporary workaround rather than a robust editing mechanism.
Future work follows tracks corresponding to these limitations. First, Longitudinal and organizational studies should examine the quality improvement in the division of labor among designers, PMs, and engineers, with attention to reviewing workflows and accountability. Second, advances in compliance by construction can integrate global tokens and accessibility rules so that violations are flagged — and fixable — directly in the semantic layer. In parallel, technical extensions are needed to address scalability and model reliability. Supporting multi-component, large-scale architectures will be essential for enabling real-world, long-term use in large and evolving UI codebases. Closed-loop learning would treat accepted edits as training signals, allowing recommender models to improve fidelity and reduce correction effort over time. Finally, Cross-modal semantics could extend beyond text to semantics inferred from sketches, wireframes, or component trees, widening on-ramps into the structured layer while maintaining traceability.

### v2

@@ 5.2. System Overview
As shown in Figure 3, users first specify semantics in the semantic structure view across the four-level framework (top-left); the system then generates the corresponding UI from these semantics and renders it (top-right). This pairing of semantic input and visual feedback anchors user intent to concrete outcomes, reducing ambiguity. Users can switch to the relation view (bottom-left), which visualizes interdependencies among semantics. From either view, clicking any semantic attribute opens the semantic detail panel for in-depth inspection of that attribute and its relationships.
@@ 5.4. Augmented Semantics Analysis
$5.$ Semantic Relationship Analysis
As shown in Figure 6, the relationship analysis renders a relation graph between semantic attributes. The graph represents semantics as nodes and their relationships as edges, with three relationship types: Values Match Well [Uncaptioned image] in green, Values Conflict [Uncaptioned image] in orange, and Needs Value [Uncaptioned image] in blue dashed lines. This graph serves dual purposes: guiding input decisions before generation and explaining implementation patterns after generation.
@@ 5.6. Iterative Refinement Support
First, the system preserves the semantic state across iterations and utilizes it in every generation. Second, to support granular, section-specific modifications, we designed a targeted regeneration strategy. While standard chat interfaces rely on the model to implicitly infer changes from a stream of conversation history, our system explicitly specifies the semantic diff (comparing previous semantics and identifying which semantic value changed and how) in the prompt. Given that the model regenerates the entire component, by providing previously generated UI code as a reference, the system instructs the model to apply changes to the corresponding modified semantics while preserving the existing structure. This strategy enables a controlled, section-specific update, preventing unintended side effects even when regenerating the full component.
@@ 6.1. Participants and Study Design
We recruited UI/UX practitioners through online design communities, resulting in 14 participants (8 UI/UX designers, 3 software engineers, 3 product managers) with varying levels of experience in UI generation tools (6 beginners, 5 intermediate users, 3 experts). Participants had an average of 4 years of professional experience (SD = 2.48, range: 1-10 years). Participants received compensation equivalent to US\$20 upon completing the session.
We employed a within-subjects design where each participant used both systems: our semantic-based interface and a baseline chat-based system with v0’s prompt enhancement. Both systems used an identical UI generation model (v0-1.5-md) to ensure that differences stemmed from the interface design rather than generation capabilities.
@@
  • Task 2 (30 mins): Open-ended screen generation with both systems, each with a different scenario.
@@
For each task, participants first generated an initial UI, then were encouraged to refine the generated UI based on their own design goals. To ensure sufficient engagement with the iterative capabilities of both systems, participants were instructed to perform a minimum of three iteration cycles per task. Beyond this requirement, they were free to continue iterating until they reached a state of subjective satisfaction. This structure allowed us to observe both the gulfs and the amplification problems.
@@ 6.3.1. Quantitative Results
Given the paired design and Likert-scale nature of our data, we employed Wilcoxon signed-rank tests (Wilcoxon, 1945) to analyze differences between systems. The tests revealed statistically significant differences across all measured dimensions (*$p<.05$, **$p<.01$) with large effect sizes ($r=Z/\sqrt{N}>.5$). Figure 7 illustrates these differences through distinct response patterns. The baseline chat interface produced polarized outcomes—users either succeeded or struggled significantly, with little middle ground. In contrast, our semantic system supported more consistent outcomes, with responses clustering toward positive ratings.
@@
Participants found our system significantly more effective for expressing design intent. They rated Intent Expressiveness substantially higher ($p$ = .008, $r$ = 0.79), with semantic system ratings (M = 5.93) notably exceeding the baseline (M = 4.64). The diverging response patterns in Figure 7 show that while baseline users frequently selected neutral or negative ratings, semantic system users concentrated on positive ratings (6-7 on the Likert scale). Intent Reflection—how accurately users felt the system understood their intentions—also showed significant improvement ($p$ = .010, $r$ = 0.71; Semantic: M = 5.79, Baseline: M = 4.36). The Think through metric demonstrated the largest effect ($p$ = .002, $r$ = 0.87), jumping from M = 3.93 in the baseline to M = 6.14 in our system, suggesting that participants perceived the structured semantic input as helpful for systematically considering design decisions rather than struggling with prompt formulation.
@@ 6.3.2. Qualitative Results
Regarding parsing accuracy, twelve participants generally found the natural language to semantic conversion accurate and helpful. P5 characterized the parsing function as “organizing my thinking for me,” while P11 was “surprised by how appropriate the categorization was.” Eleven participants also appreciated how the system flagged empty fields in relationship analysis (Needs Value), guiding them to consider overlooked or missing details. P4 described how this feature guided decision-making: “I knew I wanted a brighter design but wasn’t sure which exact one. The relation analysis recommended colors based on my ‘Target User’ and ‘Goal,’ so I applied it immediately.” P7 noted that “empty fields and suggested value showed me what was missing from my prompt,” enabling iterative completion.
@@
Regarding semantic-to-UI adherence, participants reported that specified values were generally well-reflected. P2 noted that “since the system tells me what values to input, I could specify clearly, and it felt well-reflected,” while P10 observed that abstract specifications like “warm pastel tones” resulted in coherent choices across related elements and generated UI. However, adherence varied with input specificity—“clearly specified requests came out as expected, while vague requests produced vague results” [P14]—and design novelty, with familiar UI patterns yielding higher fidelity than novel concepts [P6]. The augmented semantics (analyzing implementation) received positive feedback for extraction accuracy. P8 reported that “the analyzed results were quite accurate” and P4 noted “interaction details I would otherwise forget,” and P5 found augmented semantics helpful for discovering implicit decisions, making it easy to map design intent to which part of the UI. For semantic relationship analysis, conflict and match value indicators were perceived as actionable and reliable in understanding generated UI. P11 remarked that the “red conflict lines” acted as an immediate feedback loop, “telling me what to fix first” without needing to guess why the UI looked wrong. P10 added that seeing the connections helped explain design choices. Card-level feedback and one-click acceptance/replacement were frequently praised for reducing evaluation effort: “I can see which intent maps to which part of the screen” [P3]. Several participants linked this transparency to greater trust and faster decisions.
@@
The most distinct advantage of the semantic system was its ability to support iterative refinement, reducing “semantic drift”—the tendency for AI models to unexpectedly alter unrelated elements during updates. In the baseline chat condition, five participants (P4, P8, P11, P13, P14) explicitly described “ripple effect” failures. P14 noted: “I asked to change the top section… but it randomly changed the button colors at the bottom too.” P8 echoed this, stating that “Chat updates often reset previous design choices.” Nine participants highlighted the semantic structure’s ability to enable “scoped edits,” allowing users to lock in specific decisions. P2 and P9 highlighted the accuracy of these targeted updates, noting the ability to “target just the node I want to change,” ensuring that the rest of the interface remained stable. The accept/replace workflow of augmented semantics and relationships was described as “clickable evolution” that sped up iteration [P5, P8]. In contrast, several noted that chat-based refinement sometimes drifted, altering unrelated visual attributes or layout despite narrow requests [P4, P8, P11, P13, P14]. Participants requested lightweight change auditing—“show me exactly what changed and why”—and simple rollback/version snapshots [P1, P8, P12]. They framed these as additive features atop an already predictable loop.
@@
Across roles, participants saw the semantic approach as especially well-suited to early product exploration, structured refinement, and cross-functional collaboration. Non-designers highlighted gains in expressing and checking intent; designers and engineers pointed to component-level semantics as helpful for handoff and consistency [P9, P11, P12, P13]. Several participants wanted a bridge to lightweight direct manipulation of elements in generated UI that writes back to semantics [P4, P7, P8, P10].
Challenges and Error Responses.
Participants identified several limitations in our system. Initial confusion around semantic category boundaries was indicated—P2, P6, and P7 reported difficulty distinguishing Design Style from Visual Mood, or Function from Information Architecture. These concerns typically subsided after one or two generation cycles as participants developed intuitions for the categories. For the semantic parsing, some participants noted that long natural-language descriptions sometimes “thinned out” when compressed into semantic slots, suggesting nuance loss in parsing. Regarding relationship analysis, P1 expressed skepticism about trustworthiness (“Where’s the evidence or source?”), while P11 observed occasional cases where the system “focuses on specific intent, not for the whole intent.” Several participants [P3, P5, P7, P11] noted that both systems excel at conventional, minimal styles but struggle with highly branded or stylistically distinctive designs. These findings indicate areas for improvement in semantic granularity, parser fidelity, and support for creative divergence.
@@ 7.2. Design Implications
Designers often want to nudge spacing, alignment, or emphasis directly on canvas without losing traceability. Provide lightweight direct manipulation—handles for spacing/alignment, sliders for radius/elevation, swatch pickers—that update the bound slot and record a semantic diff. Offer “Lock to Design System” so tweaks snap to tokenized ranges. This bridge satisfies craft instincts while maintaining explainability and repeatability; start with a small set of high-impact properties and expand after stability testing.
@@ 7.3. Limitations and Future Work
Our investigation has several limitations. First, the semantic vocabulary introduces an initial learning curve; newcomers must internalize terms and relations. Progressive disclosure, examples-as-templates, and automatic slot suggestions may alleviate this burden. Second, slotting long-form intents risks nuance compression. Dual views that co-display the original text and its structured interpretation, along with preservable “verbatim” spans for critical phrasing, can mitigate loss. Third, the breadth of the evaluation remains limited: our evaluation mainly focused on perceived control, reflecting our theoretical focus on the gulfs as cognitive gaps in the user’s mental model. Consequently, the artifact-based differences, as well as longitudinal and team-level cooperative conditions, remain to be empirically validated. Additionally, technical constraints shaped the current implementation. Our system covers a single React component, which helps reduce integration complexity but limits applicability to complex UI architectures. Also, hallucination (including syntax errors in UI code) and unreliable localized editing remain challenges: instead of modifying only the intended region, the model may regenerate or alter unaffected content. Semantic diffs helped constrain these behaviors, but they served as a partial mitigation rather than a robust editing mechanism.
Future work follows tracks corresponding to these limitations. First, longitudinal and organizational studies should examine how the semantic-based approach is utilized in actual design workflows involving complex, multi-screen products, including collaboration among designers, PMs, and engineers. Second, advances in compliance by construction can integrate global tokens and accessibility rules so that violations are flagged and fixed directly in the semantic layer. In parallel, technical extensions are needed to address scalability and model reliability. Supporting multi-component, large-scale architectures will be essential for enabling real-world, long-term use in large and evolving UI codebases. Closed-loop learning would treat accepted edits as training signals, allowing recommender models to improve fidelity and reduce correction effort over time. Finally, cross-modal semantics could extend beyond text to semantics inferred from sketches, wireframes, or component trees, widening on-ramps into the structured layer while maintaining traceability.